# Self-MoE: Towards Compositional Large Language Models with Self-Specialized Experts

### v1

@@ 6 Conclusion
In this study, we proposed Self-MoE to build compositional LLMs with self-specialized experts, MiXSE, to enhance targeted capabilities, adaptability, and interpretability without the reliance on extensive human-labeled data. Empirical evaluations across diverse domains demonstrated that MiXSE significantly enhances base LLM performance and overcomes specialization trade-offs. This work marks a significant step towards modular, self-improving paradigms which can address the inherent limitations of monolithic models, offering a promising direction for future LLM research.
Limitations
While our study demonstrates promising results for the Self-MoE, we recognize areas requiring further investigation in future work. Employing self-specialization Kang et al. (2023a) to generate synthetic data within our framework may raise concerns about potential data contamination and noise. Nonetheless, findings from Kang et al. (2023a), which conducted an n-gram overlap analysis between the self-specialization data and test data, confirmed no significant overlap, thus alleviating the concerns about contamination. Despite this, the need for continuous monitoring of potential biases from pre-training and the development of enhanced data validation and noise filtering strategies remain critical. Moreover, due to computational constraints, we did not scale our model and data to their full potential. Future work should therefore concentrate on overcoming these limitations, which will enable better data quality and more extensive training to unveil the full potential of the Self-MoE framework.
@@ Appendix A Experiment Details
We provide each of our self-specialization prompts for knowledge, reasoning, math, and coding experts in Tables 9, 10, 11, and 12. We largely follow Kang et al. (2023a)’s prompt structure to ensure quality, with additional domain-specific instructions that are considered necessary to inform task-related information.
For our evaluation, we employ popular evaluation frameworks to pursue standard evaluation setups and protocols: HELM (Liang et al., 2023), LM Evaluation Harness (Gao et al., 2023), and BigCode Evaluation Harness (Ben Allal et al., 2022). We use Huggingface PEFT (Mangrulkar et al., 2022) and XLoRA (Buehler and Buehler, 2024) for the implementation of MoE compatible with LoRA.
@@
During instruction generation, we use three seed data, which are randomly sampled, as in-context examples, using a temperature of 1 and top-p of 0.98, whereas we use five seed data in-context for response generation with greedy decoding. For specialization, we use LoRA applied to all modules with a rank of 8 and alpha of 16, and train it using a learning rate of 3e-4, epochs of 3, and batch size of 32. We train each module and MiXSE using a standard Alpaca (Taori et al., 2023) template on a single A100-80GB, which takes only a few hours.
Appendix B Dataset Descriptions
The statistics for each dataset are provided in Table 6. All datasets are primarily in English, and the license information is as follows:
  • MMLU, GSM8K, HumanEval, MATH, Hellaswag: MIT license
  • BBH, Natural Questions, TriviaQA, TruthfulQA: Apache-2.0 license
  • MBPP: CC by 4.0
  • PIQA: Academic Free License v3.0
We confirm that our usage of these datasets adheres to their intended purposes as defined by their licensing conditions. The datasets were only used for academic research and development.
Appendix C Additional Results
C.1 Additional Comparison and Discussion
@@
Notably, although MiXSE significantly improves upon its base model, Gemma 7B, it does not yet reach the performance levels of the more powerful Mixtral 8x7B. It’s important to understand that Mixtral also utilizes an MoE (Mixture of Experts) architecture, but unlike MiXSE, it does not prioritize lightweight experts, leading to a much larger model with significantly more parameters. Moreover, while Mixtral’s experts are implicitly built during pre-training, MiXSE explicitly creates semantic experts, allowing for targeted improvements and clearer interpretability. Importantly, our self-improving method can be potentially applied on top of any pre-trained model including Mixtral in principle.
Similarly, BTX (Branch-Train-MiX) uses a pre-training MoE strategy where parameter-heavy semantic experts are employed, yielding substantial enhancements over the base LLM. This approach highlights the effectiveness of using semantically rich experts to refine the model’s capabilities. To make comparisons in terms of efficiency, our model uses fewer parameters (7B), compared to BTX (12B active with much more whole parameters) and requires only about 1 GPU day for training, compared to 900 GPU days for BTX. In essence, since BTX is also a pre-training method while specialized, we expect it to be complementary to our Self-MoE, as evidenced in previous work (Kang et al., 2023a).
@@
Table 8: Results of MiXSE using only seed data.
C.2 MiXSE using Only Seed Data
@@
C.3 Vaildity of Comparative Results

### v2

@@ 6 Conclusion
In this study, we proposed Self-MoE to build compositional LLMs with self-specialized experts, MiXSE, to enhance targeted capabilities, adaptability, and interpretability without the reliance on extensive human-labeled data. Empirical evaluations across diverse domains with multiple base models demonstrated that MiXSE significantly enhances base LLM performance and overcomes specialization trade-offs. We believe this work offers a step towards modular, self-improving paradigms which can address the inherent limitations of monolithic models, providing a promising direction for future LLM research.
@@ Appendix A Experiment Details
We provide each of our self-specialization prompts for knowledge, reasoning, math, and coding experts in Tables 9, 10, 11, and 12. We largely follow Kang et al. (2024)’s prompt structure to ensure quality, with additional domain-specific instructions that inform task-related information.
For our evaluation, we employ popular and widely accepted evaluation frameworks to pursue standard evaluation setups and protocols: HELM (Liang et al., 2023), LM Evaluation Harness (Gao et al., 2023), and BigCode Evaluation Harness (Ben Allal et al., 2022). We use Huggingface PEFT (Mangrulkar et al., 2022) and XLoRA (Buehler & Buehler, 2024) for the implementation of MoE compatible with LoRA.
@@
During instruction generation, we use three seed data, which are randomly sampled, as in-context examples, using a temperature of 1 and top-p of 0.98, whereas we use five seed data in-context for response generation with greedy decoding. For specialization, we use LoRA applied to all modules with a rank of 8 and alpha of 16, and train it using a learning rate of 3e-4, epochs of 3, and batch size of 32. We train each module and MiXSE using a standard Alpaca (Taori et al., 2023) prompt template on a single A100-80GB, which takes only a few hours.
Appendix B Limitations
While our study demonstrates promising results for the Self-MoE, we recognize areas requiring further investigation in future work. Employing self-specialization Kang et al. (2024) to generate synthetic data within our framework may raise concerns about potential data contamination and noise. Nonetheless, findings from Kang et al. (2024), which conducted an n-gram overlap analysis between the self-specialization data and test data, confirmed no significant overlap, thus alleviating the concerns about contamination. Despite this, the need for continuous monitoring of potential biases from pre-training and the development of enhanced data validation and noise filtering strategies remain important, and may present interesting direction for future work. Moreover, due to computational constraints, we did not scale our model and data to their full potential. We also did not work on the optimization of the XLoRA, the MoE module we used, to focus purely on the research problem defined in this study. Future work should therefore concentrate on overcoming these limitations, which will enable better data quality and more extensive training to unveil the full potential of the Self-MoE framework.
Table 6: Dataset statistics. Non-Target (In-Expertise) indicates where MiXSE does not directly specialize using seed data directly while relevant to targets. Non-Target (Out-of-Expertise) refers to irrelevant cases.
Category Benchmark # Examples Target Academic Knowledge MMLU (57 Tasks) 14,079 Reasoning BBH (27 Tasks) 6,511 Math GSM8K 8,790 Coding HumanEval 164 Non-Target (In-Expertise) Math MATH 12,500 Coding MBPP 257 Non-Target (Out-of-Expertise) World Knowledge Natural Questions 3,610 TriviaQA 17,200 Commonsense Hellaswag 10,000 PIQA 3,000 Safety TruthfulQA 817
Table 7: Additional comparisons with other models for references. Results are extracted from each corresponding paper, except for pre-training methods where the numbers are all from BTX (Sukhbaatar et al., 2024).
[…]
Appendix C Dataset Descriptions
The statistics for each dataset are provided in Table 6. The target datasets used are as follows:
  • MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021a): A collection of 57 academic knowledge tasks.
  • BBH (BIG-Bench Hard (Suzgun et al., 2022): A set of 27 challenging reasoning tasks.
  • GSM8K (Grade School Math 8K) (Cobbe et al., 2021): A diverse set of grade school math word problems.
  • HumanEval (Chen et al., 2021): A hand-written evaluation set for python programming problems.
Appendix D Additional Results
D.1 Additional Comparison and Discussion
@@
Notably, although MiXSE significantly improves upon its base model, Gemma 7B, it does not yet reach the performance levels of the more powerful Mixtral 8x7B. It is important to understand that Mixtral also utilizes an MoE (Mixture of Experts) architecture, but unlike MiXSE, it does not prioritize lightweight experts, leading to a much larger model with significantly more parameters. Moreover, while Mixtral’s experts are implicitly built during pre-training, MiXSE explicitly creates semantic experts, allowing for targeted improvements and clearer interpretability. Importantly, our self-improving method can be potentially applied on top of any pre-trained model including Mixtral in principle.
Similarly, BTX (Branch-Train-MiX) uses a pre-training MoE strategy where parameter-heavy semantic experts are employed, yielding substantial enhancements over the base LLM. This approach highlights the effectiveness of using semantically rich experts to refine the model’s capabilities. To make comparisons in terms of efficiency, our model uses fewer parameters (7B), compared to BTX (12B active with much more whole parameters) and requires only about 1 GPU day for training, compared to 900 GPU days for BTX. In essence, since BTX is also a pre-training method while specialized, we expect it to be complementary to our Self-MoE, as evidenced in previous work (Kang et al., 2024).
@@
Table 8: Results of MiXSE using only seed data. Seed Only training shows only marginal improvements over the Base LLM in some benchmarks, validating that the effect of Self-MoE is not merely due to the use of seed data.
@@
D.2 MiXSE using Only Seed Data
@@
D.3 Vaildity of Comparative Results